OpenStudy (anonymous):

multiply (x+y)^2/3 x 24/2(x+y)

OpenStudy (anonymous):

12X(x+y)^(-1/2)

OpenStudy (radar):

\[12\sqrt[3]{(x+y)^{5}}=12(x+y)\sqrt[3]{(x+y)^{2}}\]